OpenStudy (anonymous):

What is the solution of logv(x + 5)0.001 = -3 ?

OpenStudy (anonymous):

\[\log _{x+5}0.001=-3\]

OpenStudy (anonymous):

(x+5)^-3 = 0.001 solve for x. that's it

OpenStudy (anonymous):

so x=-5?

OpenStudy (anonymous):

=5 not -5

OpenStudy (anonymous):

RHS 0.001 = 10^-3